# I Don’t Know You, But I Can Catch You: Real-Time Defense against Diverse Adversarial Patches for Object Detectors

### v1

@@ Abstract.
In this paper, we propose NutNet, an innovative model for detecting adversarial patches, with high generalization, robustness and efficiency. With experiments for six detectors including YOLOv2-v4, SSD, Faster RCNN and DETR on both digital and physical domains, the results show that our proposed method can effectively defend against both the HA and AA, with only 0.4% sacrifice of the clean performance. We compare NutNet with four baseline defense methods for detectors, and our method exhibits an average defense performance that is over 2.4 times and 4.7 times higher than existing approaches for HA and AA, respectively. In addition, NutNet only increases the inference time by 8%, which can meet the real-time requirements of the detection systems. Demos of NutNet are available at: https://sites.google.com/view/nutnet.
@@ 1. Introduction
Currently, there are also some defense techniques (Xu et al., 2023; Liu et al., 2022; Chiang et al., 2021) proposed for object detection models. These typically depend on an external model for adversarial patch detection, yet these approaches still exhibit limitations in (L1) generalization, (L2) robustness and (L3) efficiency. (L1) Existing methods have limited generalization. On one hand, most of them focus on detecting HA patches, lacking validity proofs against AA patches. On the other hand, their defense models rely on training datasets containing pre-generated patches and may perform poorly on unseen patches. (L2) These methods may raise concerns about the robustness. Beyond the object detection model, they introduce a defense model to detect or filter adversarial patches, which may also be vulnerable to adversarial examples. Therefore, utilizing adversarial samples to carry out adaptive attacks against both the defense model and object detector raises concerns about the robustness of the defense model. (L3) Some methods present efficiency challenges for object detection systems since their introduced external defense models are not sufficiently lightweight (e.g., segmentation models). However, object detection models require high real-time performance since they are commonly utilized in fields such as autonomous driving or intelligent surveillance.
@@
Image-splitting amplifies the distinction between adversarial and normal inputs in feature space. In object detection scenarios, adversarial patches are typically a small part of the image, causing significant feature overlap with normal regions. To mitigate this, we divide the image into blocks, treating each as a separate input for the autoencoder. Ideally, when the block size matches the patch size, there should be minimal feature overlap between blocks with and without patches. We determine suitable block sizes considering the distribution of patch sizes observed in previous physical adversarial attacks. Destructive training further limits the model’s capability to reconstruct adversarial patches, based on the decreased feature overlap between patched and normal images. We do that by deliberately impairing the ability of the autoencoder to generate anything beyond normal samples. The decoder, with its limited generative ability, functions like a “nut”, creating a stable and self-locking system with the encoder (“screw”), hence name it NutNet. Finally, we adhere to the “less is more” design principle, introducing minimal model parameters in crafting the auto-encoder to address (L3). Our defense model, utilizing only $5k$ parameters, is significantly smaller than the segmentation models used in existing works (1M-parameter for SAC (Liu et al., 2022), APM (Chiang et al., 2021) based on U-Net, and over 25M-parameter for PatchZero (Xu et al., 2023) based on PSPNet).
We conduct extensive experiments in both the digital and physical world to validate the effectiveness of NutNet in defending six detection models, i.e., one-stage detectors (YOLOv2-v4 (Redmon and Farhadi, 2016; Redmon, Joseph and Farhadi, Ali, 2018; Bochkovskiy et al., 2020), SSD (Liu et al., 2016)), two-stage detector (Faster RCNN (Ren et al., 2016)) and transformer-based detector (DETR (Carion et al., 2020)), on four datasets, i.e., COCO, INRIA, KITTI, and Apricot. The experimental results demonstrate the effectiveness of NutNet in mitigating adversarial patches for both HA and AA. Note that defenses against HA aim to improve the $\text{AP}_{0.5}$ while defenses against AA aim to reduce the $\text{AP}_{0.5}$. Specifically, NutNet reduces $\text{AP}_{0.5}$(Average Precision evaluated with an IoU threshold of 0.5) of AA from 80.1% to 2.0%, and improves $\text{AP}_{0.5}$ of HA from 49.3% to 72.6%, outperforming existing works over 4.7 times and 2.4 times, respectively, with only a 0.4% loss of clean data accuracy. In addition, NutNet incurs very little overhead to the model detection, with only an 8% increase in the inference time, less than most existing defenses we compare. We also verify NutNet with the adaptive attack and find NutNet is still robust due to the contradiction of optimization between bypassing NutNet and achieving the adversarial attack.
@@ 2.3. Adversarial Defense
Adversarial defenses can be categorized into online defense and offline defense based on their deployment time. Offline defense refers to enhancing the robustness of the model during training, without making any modifications during the prediction stage. Offline defenses typically include adversarial training (Goodfellow et al., 2015; Tramèr et al., 2020; Madry et al., 2019; Wong et al., 2020), defensive distillation (Papernot et al., 2016), etc. Online defense, on the other hand, does not modify the model but process images during prediction, using techniques such as denoising (Gu and Rigazio, 2015; Meng and Chen, 2017a; Samangouei et al., 2018), compression (Das et al., 2017; Dziugaite et al., 2016) and smoothing (Guo et al., 2018; Naseer et al., 2018). These defenses are generally effective only against adversarial perturbations. For adversarial patches against classifiers, some explanation-based methods like (Hayes, 2018; Chen et al., 2021; Zhang et al., 2020) use heatmaps or saliency maps to locate the patches and then apply masking or suppression. However, explanation-based methods are unsuitable for object detectors, as the detector itself cannot identify the attack type. Therefore, defenses for adversarial patches on detectors typically rely on sliding masking (Xiang et al., 2022b) or external models (Xiang and Mittal, 2021; Liu et al., 2022; Xu et al., 2023) to mitigate the patches.
@@ 3.1. Threat Model
Attacker’s goal and capability. The attacker aims to fool a real-time object detector by adversarial patches. It is assumed that the attacker has full knowledge of the victim detector, including its parameters and structures, and can generate effective patches using white-box methods. The attacker can manipulate the attack type (HA or AA), size, shape, position, and number of patches used, without prior knowledge by the defenders. Furthermore, the attacker might be aware of the defense deployment, enabling him to launch adaptive attacks against the defense method.
Defender’s goal and capability. The defender aims to make the detector output the correct prediction without being affected by adversarial attacks. However, the defender is unaware of whether the detector’s input is a clean image or an image with an adversarial patch. Such a threat model maximizes the adversary’s capabilities and aligns with the real-world scenario faced by the defenders. The defender employs an online defense approach to process the images instead of employing an offline defense approach to modify the victim model since an offline defense cannot adapt to different types of adversarial attacks and may not handle unseen attack scenarios.
@@ 3.2. Motivation and Defense Intuition
To achieve this, we need to address two questions: how to make the defense model small and how to improve its performance and robustness. Intuitively, the model can perform well with few parameters when it doesn’t need to learn excessive and complex knowledge. As a solution, we deconstruct the intricate functionality of existing defense models, encompassing tasks such as detecting, segmenting, locating, and masking adversarial patches. We preserve only the detection function within the model, employing non-model methods for other functions. This approach leads to a substantial reduction in the model’s size, making it lightweight. To enhance the lightweight model’s ability to detect adversarial patches from input images, we further analyze the differences between adversarial patches and clean images, finding adversarial patches always differ from clean images in semantics, textures, or colors. Therefore, treating adversarial patches as OOD samples for detection, coupled with widening the gap between the distributions of adversarial patches and clean images, provides a potential solution to improve the performance and robustness.
@@ 3.3. Overview of NutNet
The defense pipeline. NutNet follows a five-step process during the defense phase. Firstly, the input image is split into small non-overlapping blocks. These blocks are subsequently input into NutNet, serving as a distribution extractor, responsible for reconstructing the blocks. Through an analysis of the reconstruction error, a coarse-grained block-wise mask and a fine-grained pixel-wise mask are generated. These masks are then combined to create a more precise final mask. Subsequently, the patch region is excised from the image based on the final mask, preparing the modified image for detection by the detector.
@@ 4.1. NutNet
We assume that clean images and adversarial patches originate from different distributions. Therefore, we model these two distributions and employ the concept of out-of-distribution detection to identify adversarial patches. Specifically, we can train a distribution extractor to allow only images sampled from the clean distribution to pass through, effectively blocking images from other distributions, thus we can defend against adversarial patches. We name this distribution extractor NutNet because, like a nut restricting the passage of screws that do not meet its constraints, NutNet imposes strong constraints on its inputs. Next, we will delve into NutNet from two perspectives: model architecture and training methodology.
To differentiate the clean sample patterns and adversarial patch patterns based on the data distribution, we first need to establish a model for the data distribution and then measure the distance of a specific sample from this distribution. Therefore, we exploit a reconstruction-based autoencoder to learn the representation of the clean data distribution since we have access to the clean training dataset. We achieve this by training it to accurately construct the clean samples but fail to reconstruct non-clean samples effectively. Following this, we view reconstruction error as a form of distance measurement between the input sample and the clean data distribution.
@@
The Image-splitting Strategy. For NutNet to precisely extract the distribution of input images, we need to constrain its input dimensions. Considering the stealthiness of attacks, adversaries often limit the size of patches. Consequently, in a given input image, an adversarial patch typically occupies only a small fraction. If the autoencoder’s input is a complete image, it becomes challenging to meet our requirements for distribution extraction. This is because the majority of regions in such an image belong to the distribution of clean images, which is the distribution we aim to accurately reconstruct. This complexity makes it difficult for the autoencoder to identify the distribution of small areas containing adversarial patches in the image.
@@
Therefore, we propose the Image-splitting strategy to divide the image into smaller, non-overlapping blocks for extraction using NutNet. The Image-splitting addresses the issue mentioned earlier, where adversarial patches occupy only a small fraction. With appropriate block size, some blocks will contain the majority of adversarial patches, while the rest contain almost clean images. This magnifies the contrast between clean and anomalous blocks, making it easier for the autoencoder to learn how to differentiate their distributions. Additionally, compared to processing the entire input image, an autoencoder that handles only small image blocks can have fewer parameters and a faster inference speed. The fewer parameters simplify the autoencoder’s feature space, limiting the potential attack surface exposed to adaptive attacks. The faster inference speed minimizes the additional overhead introduced by defense, facilitating real-time detection for the object detector.
@@
Finally, we need to choose an appropriate size for Image-splitting. While adversarial patches cannot be too large for stealthiness, they also cannot be too small for effectiveness. We aim to select an appropriate size that covers the smallest yet effective adversarial patches. In our investigation of existing adversarial attacks, HA patches for hiding a person are usually effective within a range of a few meters, while AA patches for causing a stop sign to appear can be effective from distances exceeding 20 meters. Analyzing the dimensions of the adversarial patches used, we find that the minimum height of effective patches typically needs to be around $1/30$ of the image height. Therefore, we can choose this size for splitting. Taking the input size of $416\times 416$ for YOLO detectors as an example, we select a size of $13\times 13$ $(416=13\times 32)$ for splitting the image. In addition, we can also use a size of $26\times 26$, or $52\times 52$. The smaller the block size, the easier it is for $\mathcal{E}$ to filter out patches, but dividing an image into too small blocks may also lead to false positives. The larger the block size, the more features of the image are retained, making it less likely to generate false positives, but the filtering ability for small patches may also decrease. Therefore, we can choose different block sizes for practical applications according to different security requirements. The impact of different block sizes on detection accuracy is evaluated in Section 5.7.
@@
Formally, we use $P_{c}$ to represent the distribution of clean images (i.e., $x\sim P_{c}$). The autoencoder $\mathcal{E}$ encodes the input images with a convolutional network and then decodes to output the reconstruction of original images with another convolutional network. Let $Dist(\cdot,\cdot)$ be the distance (i.e., reconstruction error) of two images, and we hope that $\mathcal{E}$ achieves $Dist(x,\mathcal{E}(x))\ll Dist(x^{\prime},\mathcal{E}(x^{\prime}))$, where $x\sim P_{c},x^{\prime}\nsim P_{c}$. Therefore, the objective function to train $\mathcal{E}$ is as follows:
@@ 5.1. Experimental Setup
Object detectors. Due to the decoupling between NutNet and the object detectors, it is easy to evaluate NutNet on different object detectors. We select YOLOv2-v4 (Redmon and Farhadi, 2016; Redmon, Joseph and Farhadi, Ali, 2018; Bochkovskiy et al., 2020), SSD (Liu et al., 2016), Faster RCNN (Ren et al., 2016) and DETR (Carion et al., 2020) as the victim object detectors, which include traditional convolutional one-stage, two-stage detection architectures and the detection transformer. The backbones of YOLOv2-4 and SSD are Darknet-19, Darknet-53, CSPDarknet-53 and VGG-16 respectively, while the backbones of Faster RCNN and DETR are ResNet-50. SSD is trained on the VOC (PASCAL Visual Object Classes) (Mark et al., 2010) dataset while the other models are trained on COCO (Microsoft Common Objects in Context) (Lin et al., 2015) dataset.
@@
Patch sizes. For adversarial patch attacks in the digital world, we set different sizes for different patch types. For the targeted HA patches, we scale each patch to one-fifth of the length of the diagonal of the corresponding bounding box of the object and paste them over target objects, following the setup of (Thys et al., 2019). To enhance the attack success rate of patches with lower efficacy at this size, we further increase their dimensions by 25%, resulting in an improved attack success rate. Comprehensive details regarding the size of the adversarial patches will be provided in Appendix B.2. In the case of untargeted HA patches, we scale each patch following the specifications outlined in (Lee and Kolter, 2019) and affix them to the corners of the image to minimize interference with other objects. As for the AA patches, we randomly adjust the size and position them within the image, simulating the capturing of these objects at varying distances. For physical-world adversarial patch attacks, we print the patches at a size of 40cm by 40cm, aligning with the setup detailed in (Thys et al., 2019). The natural patches are printed at a size of 50cm by 50cm, as natural patches do not achieve the same level of attack effectiveness as regular patches of identical size.

### v2

@@ Abstract.
In this paper, we propose NutNet, an innovative model for detecting adversarial patches, with high generalization, robustness and efficiency. With experiments for six detectors including YOLOv2-v4, SSD, Faster RCNN and DETR on both digital and physical domains, the results show that our proposed method can effectively defend against both the HA and AA, with only 0.4% sacrifice of the clean performance. We compare NutNet with four baseline defense methods for detectors, and our method exhibits an average defense performance that is over 2.4 times and 4.7 times higher than existing approaches for HA and AA, respectively. In addition, NutNet only increases the inference time by 8%, which can meet the real-time requirements of the detection systems. Demos and the full paper of NutNet are available at: https://sites.google.com/view/nutnet.
@@ 1. Introduction
Currently, there are also some defense techniques (Xu et al., 2023; Liu et al., 2022; Chiang et al., 2021) proposed for object detection models. These typically depend on an external model for adversarial patch detection, yet these approaches still exhibit limitations in (L1) generalization, (L2) robustness and (L3) efficiency. (L1) Existing methods have limited generalization. On one hand, most of them focus on detecting HA patches, lacking validity proofs against AA patches. On the other hand, their defense models rely on training datasets containing pre-generated patches and may perform poorly on unseen patches. (L2) These methods may raise concerns about the robustness. Beyond the object detection model, they introduce a defense model to detect or filter adversarial patches, which may also be vulnerable to adversarial examples. Therefore, utilizing adversarial examples to carry out adaptive attacks against both the defense model and object detector raises concerns about the robustness of the defense model. (L3) Some methods present efficiency challenges for object detection systems since their introduced external defense models are not sufficiently lightweight (e.g., segmentation models). However, object detection models require high real-time performance since they are commonly utilized in fields such as autonomous driving or intelligent surveillance.
@@
Image-splitting amplifies the distinction between adversarial and normal inputs in feature space. In object detection scenarios, adversarial patches typically occupy only a small portion of the image, making them difficult to differentiate. To mitigate this, we divide the image into blocks, treating each as a separate input for the autoencoder. Ideally, when the block size matches the patch size, there should be minimal feature overlap between blocks with and without patches. We determine suitable block sizes considering the distribution of patch sizes observed in previous physical adversarial attacks. Destructive training further limits the model’s capability to reconstruct adversarial patches on the basis of decreased feature overlap between patched and normal images. We do that by deliberately impairing the ability of the autoencoder to generate anything beyond normal samples. The decoder, with its limited generative ability, functions like a “nut”, creating a stable and self-locking system with the encoder (“screw”), hence we name it NutNet. Finally, we adhere to the “less is more” design principle, introducing minimal model parameters in crafting the autoencoder to address (L3). Our defense model, utilizing only $5k$ parameters, is significantly smaller than the segmentation models used in existing works (1M-parameter for SAC (Liu et al., 2022), APM (Chiang et al., 2021) based on U-Net, and over 25M-parameter for PatchZero (Xu et al., 2023) based on PSPNet).
We conduct extensive experiments in both the digital and physical world to validate the effectiveness of NutNet in defending six detection models, i.e., one-stage detectors (YOLOv2-v4 (Redmon and Farhadi, 2016; Redmon, Joseph and Farhadi, Ali, 2018; Bochkovskiy et al., 2020), SSD (Liu et al., 2016)), two-stage detector (Faster RCNN (Ren et al., 2016)) and transformer-based detector (DETR (Carion et al., 2020)), on four datasets, i.e., COCO, INRIA, KITTI, and Apricot. The experimental results demonstrate the effectiveness of NutNet in mitigating adversarial patches for both HA and AA. Note that defenses against HA aim to improve the $\text{AP}_{0.5}$ (Average Precision evaluated with an IoU threshold of 0.5) while defenses against AA aim to reduce the $\text{AP}_{0.5}$. Specifically, NutNet reduces $\text{AP}_{0.5}$ of AA from 80.1% to 2.0%, and improves $\text{AP}_{0.5}$ of HA from 49.3% to 72.6%, outperforming existing works over 4.7 times and 2.4 times, respectively, with only a 0.4% loss of clean data accuracy. In addition, NutNet incurs very little overhead to the model detection, with only an 8% increase in the inference time, less than most existing defenses we compare. We also verify NutNet with the adaptive attack and find NutNet is still robust due to the contradiction of optimization between bypassing NutNet and achieving the adversarial attack.
@@ 2.3. Adversarial Defense
Adversarial defenses can be categorized into online defense and offline defense based on their deployment time. Offline defense refers to enhancing the robustness of the model during training, without making any modifications during the prediction stage. Offline defenses typically include adversarial training (Goodfellow et al., 2015; Tramèr et al., 2020; Madry et al., 2019; Wong et al., 2020), defensive distillation (Papernot et al., 2016), etc. Online defense, on the other hand, does not modify the model but process images during prediction, using techniques such as denoising (Gu and Rigazio, 2015; Meng and Chen, 2017a; Samangouei et al., 2018), compression (Das et al., 2017; Dziugaite et al., 2016) and smoothing (Guo et al., 2018; Naseer et al., 2018). These defenses are generally effective only against adversarial perturbations. For adversarial patches against classifiers, some explanation-based methods like (Hayes, 2018; Chen et al., 2021; Zhang et al., 2020) use heatmaps or saliency maps to locate the patches and then apply masking or suppression. However, the output of a detector is more complex than that of a classifier, and different types of attacks affect the detector’s output in various ways. Consequently, it is challenging to directly apply explanation-based methods designed for classifiers to detectors. Therefore, defenses for adversarial patches on detectors typically rely on sliding masking (Xiang et al., 2022b) or external models (Xiang and Mittal, 2021; Liu et al., 2022; Xu et al., 2023) to mitigate the patches.
@@ 3.1. Threat Model
Attacker’s goal and capability. The attacker aims to fool a real-time object detector by adversarial patches. It is assumed that the attacker has full knowledge of the victim detector, including its parameters and structures, and can generate effective patches using white-box methods. The attacker can manipulate the attack type (HA or AA), size, shape, position, and number of patches used, without prior knowledge by the defenders. Furthermore, the attacker might be aware of the defense deployment including the algorithm and defense model, enabling him to launch adaptive attacks against the defense method.
Defender’s goal and capability. The defender aims to make the detector output the correct prediction without being affected by adversarial attacks. However, the defender is unaware of whether the detector’s input is a clean image or an image with an adversarial patch. Such a threat model maximizes the adversary’s capabilities and aligns with the real-world scenario faced by the defenders. The defender employs an online defense approach to process the images instead of employing an offline defense approach to modify the victim model.
@@ 3.2. Motivation and Defense Intuition
To achieve this, we need to address two questions: how to make the defense model small and how to improve its performance and robustness. Intuitively, a model can perform well with few parameters when it doesn’t need to learn excessive and complex knowledge. As a solution, we deconstruct the intricate functionality of existing defense models, encompassing tasks such as detecting, segmenting, locating, and masking adversarial patches. We preserve only the detection function within the model, employing non-model methods for other functions. This approach leads to a substantial reduction in the model’s size, making it lightweight. To enhance the lightweight model’s ability to detect adversarial patches from input images, we further analyze the differences between adversarial patches and clean images, finding adversarial patches always differ from clean images in semantics, textures, or colors. Therefore, treating adversarial patches as OOD samples for detection, coupled with widening the gap between the distributions of adversarial patches and clean images, provides a potential solution to improve the performance and robustness.
@@ 3.3. Overview of NutNet
The defense pipeline. NutNet contains the following processes during the defense phase. Firstly, the input image is split into small non-overlapping blocks, which are subsequently fed into NutNet. NutNet serves as a distribution extractor, which is responsible for reconstructing the blocks. Through an analysis of the reconstruction error, a coarse-grained block-wise mask and a fine-grained pixel-wise mask are generated. These masks are then combined to create a more precise final mask. Subsequently, the patch region is excised from the image based on the final mask, preparing the modified image for detection by the detector.
@@ 4.1. NutNet
We assume that clean images and adversarial patches originate from different distributions. Therefore, we model these two distributions and employ the concept of out-of-distribution detection to identify adversarial patches. Specifically, we can train a distribution extractor to allow only images sampled from the clean distribution to pass through, effectively blocking images from other distributions, thus we can defend against adversarial patches. We name this distribution extractor NutNet because, like a nut restricting the passage of screws that do not meet its constraints, NutNet imposes strong constraints on its inputs.
To differentiate the clean sample patterns and adversarial patch patterns based on the data distribution, we first need to establish a model for the data distribution and then measure the distance of a specific sample from this distribution. Therefore, we exploit a reconstruction-based autoencoder to learn the representation of the clean data distribution since we have access to the clean training dataset. We achieve this by training it to accurately construct the clean samples but fail to reconstruct non-clean samples effectively. Following this, we view reconstruction error as a form of distance measurement between the input sample and the clean data distribution. Next, we will delve into NutNet from two perspectives: model architecture and training methodology.
@@
The Image-splitting Strategy. For NutNet to precisely extract the distribution of input images, we need to constrain its input dimensions. Considering the stealthiness of attacks, the attackers often limit the size of patches. Consequently, in a given input image, an adversarial patch typically occupies only a small fraction. If the autoencoder’s input is a complete image, it becomes challenging to meet our requirements for distribution extraction. This is because the majority of regions in such an image belong to the distribution of clean images, which is the distribution we aim to accurately reconstruct. This complexity makes it difficult for a lightweight autoencoder to identify the distribution of small areas containing adversarial patches in the image.
@@
Therefore, we propose the Image-splitting strategy to divide the image into smaller, non-overlapping blocks for a more precise extraction. The Image-splitting addresses the issue mentioned earlier, where adversarial patches occupy only a small fraction. With appropriate block size, some blocks will contain the majority of adversarial patches, while the rest contain almost clean images. This magnifies the contrast between clean and anomalous blocks, making it easier for the autoencoder to learn how to differentiate their distributions. Additionally, compared to processing the entire input image, an autoencoder that handles only small image blocks can have fewer parameters and a faster inference speed. The fewer parameters simplify the autoencoder’s feature space, limiting the potential attack surface exposed to adaptive attacks. The faster inference speed minimizes the additional overhead introduced by defense, facilitating real-time detection for the object detector.
@@
Finally, we need to choose an appropriate size for Image-splitting. While adversarial patches cannot be too large for stealthiness, they also cannot be too small for effectiveness. We aim to select an appropriate size that covers the smallest yet effective adversarial patches. In our investigation of existing adversarial attacks, HA patches for hiding a person are usually effective within a range of a few meters, while AA patches for causing a stop sign to appear can be effective from distances exceeding 20 meters. Analyzing the dimensions of the adversarial patches used, we find that the minimum height of effective patches typically needs to be around $1/30$ of the image height. Therefore, we can choose this size for splitting. Taking the input size of $416\times 416$ for YOLO detectors as an example, we select a size of $13\times 13$ $(416=13\times 32)$ for splitting the image. In addition, we can also use a size of $26\times 26$, or $52\times 52$. The smaller the block size, the easier it is for the extractor to filter out patches, but dividing an image into too small blocks may also lead to false positives. The larger the block size, the more features of the image are retained, making it less likely to generate false positives, but the filtering ability for small patches may also decrease. Therefore, we can choose different block sizes for practical applications according to different security requirements. The impact of different block sizes on detection accuracy is evaluated in Section 5.7.
@@
Formally, we use $P_{c}$ to represent the distribution of clean images (i.e., $x\sim P_{c}$). The autoencoder $\mathcal{E}$ encodes the input images with a convolutional network and then decodes to output the reconstruction of original images with another transposed convolutional network. Let $Dist(\cdot,\cdot)$ be the distance (i.e., reconstruction error) of two images, and we hope that $\mathcal{E}$ achieves $Dist(x,\mathcal{E}(x))\ll Dist(x^{\prime},\mathcal{E}(x^{\prime}))$, where $x\sim P_{c},x^{\prime}\nsim P_{c}$. Therefore, the objective function to train $\mathcal{E}$ is as follows:
@@ 5.1. Experimental Setup
Object detectors. Due to the decoupling between NutNet and the object detectors, it is easy to evaluate NutNet on different object detectors. We select YOLOv2-v4 (Redmon and Farhadi, 2016; Redmon, Joseph and Farhadi, Ali, 2018; Bochkovskiy et al., 2020), SSD (Liu et al., 2016), Faster RCNN (Ren et al., 2016) and DETR (Carion et al., 2020) as the victim object detectors, which include traditional convolutional one-stage, two-stage detection architectures and the transformer-based architecture. The backbones of YOLOv2-4 and SSD are Darknet-19, Darknet-53, CSPDarknet-53 and VGG-16 respectively, while the backbones of Faster RCNN and DETR are ResNet-50. SSD is trained on the VOC (PASCAL Visual Object Classes) (Mark et al., 2010) dataset while the other models are trained on COCO (Microsoft Common Objects in Context) (Lin et al., 2015) dataset.
@@
Patch sizes. For adversarial patch attacks in the digital world, we set different sizes for different patch types. For the targeted HA patches, we scale each patch to one-fifth of the length of the diagonal of the corresponding bounding box of the object and paste them over target objects, following the setup of (Thys et al., 2019). To enhance the attack success rate of patches with lower efficacy at this size, we further increase their dimensions by 25%, resulting in an improved attack success rate. Comprehensive details regarding the size of the adversarial patches will be provided in Appendix B.2. In the case of untargeted HA patches, we scale each patch following the specifications outlined in (Lee and Kolter, 2019) and affix them to the corners of the image to minimize interference with other objects. As for the AA patches, we randomly adjust the size and position them within the image, simulating the capturing of these objects at varying distances. For physical-world adversarial patch attacks, we print the patches at a size of 40cm by 40cm, aligning with the setup detailed in (Thys et al., 2019). The patches with natural lookings are printed at a size of 50cm by 50cm, as those patches do not achieve the same level of attack effectiveness as regular patches of identical size.